Two point charges are separated by a distance of 14.6 cm. The charge on the one of them is 1.58 uC, and the Electric Force between the two charges if 6.15 N. Find the charge on the second particle in uC.
Answer #1

Electrostatic force is given by:

F = k*q1*q2/R^2

Given that

F = Force between the two charges = 6.15 N

q1 = charge on 1st particle = 1.58 uC = 1.58*10^-6 C

q2 = charge on 2nd particle = ?

R = distance between both charge = 14.6 cm = 0.146 m

k = 9*10^9 = electrostatic constant

So,

q2 = F*R^2/(k*q1)

Using known values:

q2 = 6.15*0.146^2/(9*10^9*1.58*10^-6)

q2 = 9.22*10^-6 C

q2 = 9.22 uC

Please Upvote.